# Robust Utility Optimization via a GAN Approach

### v1

@@ 4.3.3 Evaluation Criterion: Early Stopping and General Assessment of Robustness
Since no analytic solution is known, we have to define new metrics to asses the robustness and quality of trained models, and to perform early stopping during training. We propose the following metrics, which are all based on a reference market model. In real world settings, this reference market model is obtained by selecting a period of historical test data on which the reference parameters $\tilde{\Sigma},\tilde{\mu}$ and their respective (co)variance are estimated. For synthetic experiments, we assume that reference parameters $\tilde{\Sigma},\tilde{\mu}$ and their (co)variance are given.
@@
    We remark that the cumulative noise type is the most realistic one for real world market behaviour, since parameter uncertainty increases with time. In the subsequent evaluations of Section 4.3.4, we therefore focus on cumulative noise. In this setting, the market scenario $S^{j}$ is thus given by the SDE (12) with parameters $\Sigma^{\mathrm{noisy},j}_{t},\mu^{\mathrm{noisy},j}_{t}$ (where the Cholesky decomposition is used to compute a standard deviation matrix $\sigma$ from a variance matrix $\Sigma$).
@@ 4.3.4 Experimental Results
In this section, we report the results of a series of experiments on realistic penalized robust utility optimization as outlined earlier in Section 4.3. In these experiments, we train and compare non-robust, robust volatility and fully robust models in a 2-dimensional market with setting (AS) for the reference volatility $\tilde{\Sigma}$, setting (AD) for the reference drift $\tilde{\mu}$, and all of interest rate, step size, and initial asset value as in Section 4.1. We include proportional transaction costs $c_{\text{prop}}=1\%$, consider the three utility functions $u\in\{u_{1/2},u_{1}=\ln,u_{2}\}$ and path-dependent penalties as defined in Section 4.3.2.
@@
For the model evaluation on the test set we use the standard deviations of $0.075$ for the volatility and $0.01$ for the drift. To counteract the too optimistic view of the validation metric (19), we double these standard deviations, i.e., $0.15$ for the volatility and $0.02$ for the drift, for the early stopping validation during training. Moreover, we use $n_{\text{pool}}=1000$, $B_{\text{test}}=B_{\text{val}}=40K$ test and validation samples and $160K$ training samples. As outlined earlier in Section 4.3.3, we only report results with the cumulative noise type.
In Figure 3 (left) we show the results of a grid search over the scaling factors $\lambda_{1},\lambda_{2}\in\{0.01,0.1,0.5,1,10,100\}$, which scale the penalties $R_{1}$ and $R_{2}$ for the quadratic covariation $l_{\text{QCV}}$ and for the average relative return $l_{\text{ARR}}$, respectively, for the 3 utility functions $u_{1/2},u_{1},u_{2}$.
@@
Moreover, in Table 1 we see that for all 3 utility functions, the best combination of $\lambda_{1},\lambda_{2}$ leads to the pooled minimum expected utilities $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$, which are better (higher) than the baseline of investing everything in the risk-free bank account (given by $u(X_{0}\,e^{rT})$) and higher than the results of strategies trained in the non-robust setting, corresponding to $\lambda_{1}=\lambda_{2}=\infty$, as well as in the half-robust setting, where only the volatility is robustified, corresponding to $\lambda_{2}=\infty$ (here the maximum is always attained at $\lambda_{1}=0.01$).
@@
In particular, we see that the robust training leads to reasonable strategies, when using the appropriate penalty scaling parameters $\lambda_{1}$ and $\lambda_{2}$ for the given noisy evaluation market models and utility function. We also see that for all utility functions the optimal scaling parameters are close ($(\lambda_{1},\lambda_{2})=(0.5,10)$ for $u_{2}$, $(\lambda_{1},\lambda_{2})=(0.5,0.5)$ for $u_{1}$, $(\lambda_{1},\lambda_{2})=(1,1)$ for $u_{1/2}$). In real world settings this means that one has to first decide or understand against which type or magnitude of uncertainty one wants to robustify and then has to find appropriate penalty scaling factors for the respective setting, e.g., via grid search. Note that when comparing (minimal) expected utilities across different utility functions, it is important to keep in mind that larger utility values do not imply favorable distributions of the terminal wealth $X_{T}$, since different utility functions assign different values to the same terminal wealth (cf. Figure 2).
In Figure 4, we show test sample paths and mean $\pm$ standard deviations over all sample paths for the best strategy with utility $u_{1/2}$ evaluated against the trained market discriminator model and against one cumulative noisy market scenario. Especially in the average plots (the right panels of Figure 4), we see that the trained market and the noisy market behave differently. Nevertheless, the trained strategy works well when evaluated against the noisy market model. In particular, the average behaviour of the strategy and the resulting average portfolio value evolution are very similar in both cases, even though the market behaves differently.
@@
Similarly to before in Sections 4.1 and 4.2, FFNNs performed on par with RNNs (tested with the utility function $u_{1}$). This finding suggests that also in this more realistic and complex setup the optimal strategy is not path-dependent.
@@
In Figure 3 (right) we show the results of a grid search over scaling factors $\lambda_{1},\lambda_{2}\in\{0.01,0.5,0.1,1.,10,100\}$ for the path-dependent penalties w.r.t. the quadratic covariation $l_{\text{QCV}}$ and the average relative return $l_{\text{ARR}}$, for the utility functions $u_{1/2},u_{1},$ and $u_{2}$. In Table 2, we see that for all 3 utility functions, the best combination of $\lambda_{1},\lambda_{2}$ leads to the pooled-minimum expected utilities $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ that exceed the baseline of investing everything in the risk-free bank account (given by $u(X_{0}\,e^{rT})$). Moreover, $\max_{\lambda_{1},\lambda_{2}}M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ is also higher than the results of strategies trained in the non-robust setting, corresponding to $\lambda_{1}=\lambda_{2}=\infty$, as well as in the half-robust setting, where only the volatility is robustified, corresponding to $\lambda_{2}=\infty$ (which is always attained for $\lambda_{1}=0.01$).
In total, the evaluation against the noisy GARCH models shows that the robust model outperforms the others, when choosing the scaling factors correctly. Interestingly, the fully robust model for $u_{1}$ leads to a slightly higher value of $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ than for $u_{1/2}$. This is surprising, since $u_{1/2}$ dominates $u_{1}$ and one would expect this monotonicity to translate to the metric. [^2] In particular, using the strategy that results from the optimization w.r.t. $u_{1}$ in the evaluation with $u_{1/2}$ has to lead to the same or higher expected utility than the evaluation w.r.t. $u_{1}$. As we can see in Figure 3, the best combination of scaling factors is always attained at a boundary point of the grid. Hence, it is expected that better scaling factors could be found, when further enlarging the grid. In line with this, we expect that then also the optimal combination for $u_{1/2}$ leads to a larger value of $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ than for $u_{1}$.
@@
Figure 5: Histograms of discounted terminal portfolio value when evaluating a trained strategy against a constant baseline measure $(\tilde{\mu},\tilde{\Sigma})$ for fully robust (left) and non-robust (right) models, with utility function $u_{1/2}$ (top) and $u_{2}$ (bottom). Expected Utilities (left to right, top to bottom): 0.0164, 0.0203, 0.0153, 0.0160.
In Figure 5 we compare the distribution of discounted terminal wealth $e^{-rT}\,X_{T}$ resulting from trading with the fully robust and non-robust strategies of Table 1 $\pi^{\text{NN}}_{\text{fully rob}}$ and $\pi^{\text{NN}}_{\text{non rob}}$, respectively, obtained by (robust) portfolio optimization with the utility functions $u_{2}$ and $u_{1/2}$. We see that the trading with robust strategies leads to lower mean terminal portfolio value (PV), however, also manages to reduce the termianal PV’s standard deviation. Hence, even though there is some cost for the robustification, it also reduces the risk of the strategy. Additionally, the robust strategies have the advantage of being right-skewed (large profits are more likely to happen than large losses) compared to the non-robust ones which are not skewed (large profits and large losses are similarly likely to happen). Except for the skew, we see that changing from the non-robust (top-right) to the robust (top-left) strategy has a qualitatively similar effect as changing from a more risk-seeking (top-right) to a more risk-avers (bottom-right) utility function. Doing both (from top-right to bottom-left) increases the effect.
@@
In this section, we study in more detail the effect of small, proportional trading costs on optimal investments in both robust and non-robust utility optimization. In this paper, we so far tested how a GAN-based algorithm (cf. Section 3) can be leveraged to recover known analytical solutions to robust utility optimization (Section 4.1) and enable us to go beyond simple, friction-less settings to find robust strategies in presence of proportional trading costs and for general power utilities (Sections 4.2 and 4.3). While we saw benefits of GAN-based trading strategies in expected utilities under noisy market evaluations in Sections 4.2 and 4.3, in this Section 4.4, we investigate potential benefits of the learned trading strategies compared to other well known, asymptotically optimal investments under vanishing trading costs.
In non-robust portfolio optimization, such an analysis of optimal trading strategies in presence of (vanishingly) small trading costs has been performed in Muhle-Karbe et al. (2017). For general power utilities of the form $\tilde{u}_{p}(x)=\tfrac{x^{1-p}}{1-p}$, for $p\in(0,1)$ (note that this definition slightly differs from our previous definition in Equation 17), Muhle-Karbe et al. (2017) derive closed form approximations of asymptotically optimal trading strategies in presence of small trading costs for Black–Scholes and more general mean-reverting financial markets. In this section, we take up their discussion by comparing our GAN-based robust and non-robust strategies to the asymptotically optimal strategy derived in Muhle-Karbe et al. (2017) for Black–Scholes markets in Section 4.4.1. Then, in Section 4.4.2, we compare those strategies in one-dimensional markets consisting of a risky asset with heavy tailed log-returns.
@@ 4.4.1 Comparison in Black–Scholes Markets
In our experiments we use a similar setup as before with $p=0.5$, $c_{\text{prop}}=1\%$, $X_{0}=1$, $S_{0}=1$, $\tilde{\sigma}=35\%$, $\delta_{t}=1/65$ with $65$ steps, i.e., $T=1$, and two different regimes for the drift and risk-free rate $(\tilde{\mu}_{S},r)\in\{(4\%,1.5\%),(7\%,3\%)\}$, hence, $\tilde{\mu}\in\{5.5\%,10\%\}$. For the robust training we use the path-dependent penalties of Section 4.3.2 and grid search over $\lambda_{1},\lambda_{2}\in\{0.01,0.5,0.1,1.,10,100\}$. For validation of robust models, we consider noisy market models around the reference market, with the same setup as in the corresponding paragraph in Section 4.3.4. For the non-robust model, validation is done without noise, i.e., with the reference market with characteristics $\tilde{\mu}$ and $\tilde{\sigma}$ by approximating (13) with $B_{\text{val}}=40K$ independent validation sample paths. For each experiment, we run a range of models with various architectures and select the model that performed best in the validation for final evaluation on a separate test set. In the evaluation, we approximate the metric (18) with $n_{\text{pool}}=1000$ and $B_{\text{test}}=40K$, and the metric (13) with $B_{\text{test}}=100K$.
@@ Results
In Table 3, we show the comparison of the reference strategy $\pi^{\text{ref}}$ and the trained strategies $\pi^{\text{NN}}$ evaluated against the noisy market models and against the reference markets (with highest, i.e., best values per column in bold). For the fully robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$, we show the results of the best combination of scaling factors, given by $\lambda_{1}=10,\lambda_{2}=1$ (for $\tilde{\mu}=5.5\%$ and $r=1.5\%$), and by $\lambda_{1}=10,\lambda_{2}=10$ (for $\tilde{\mu}=10\%$ and $r=3\%$), respectively.
Table 3: Numerical results of reference strategy and trained GAN strategies evaluated against noisy market models with (18) and against the reference market with (13). Additionally, the discounted value at risk $\operatorname{VaR}_{\alpha}$ for $\alpha=5\%$ is computed in the noise-free reference markets, corresponding to the quantiles of the terminal portfolio value plotted in Figure 7.
@@
$M_{\tilde{u}_{p}}(\pi)$ $E_{\tilde{u}_{p}}(\pi,\tilde{\mu},\tilde{\sigma})$ $\operatorname{VaR}_{\alpha}$ $M_{\tilde{u}_{p}}(\pi)$ $E_{\tilde{u}_{p}}(\pi,\tilde{\mu},\tilde{\sigma})$ $\operatorname{VaR}_{\alpha}$
@@
In the non-robust case, we see that the GAN strategy $\pi^{\text{NN}}_{\text{non rob}}$ performs on par with the reference strategy $\pi^{\text{ref}}$ in the evaluation on the reference market ($E_{\tilde{u}_{p}}$ entries of Table 3). The slightly higher result for $E_{\tilde{u}_{p}}(\pi^{\text{NN}}_{\text{non rob}},\tilde{\mu},\tilde{\sigma})$ in the case of $\tilde{\mu}=10\%,r=3\%$ is most likely an effect of time discretization and the fact that the reference strategy is only asymptotically optimal for vanishing trading cost. [^3] The values for $E_{\tilde{u}_{p}}(\pi^{\text{NN}}_{\text{non rob}},\tilde{\mu},\tilde{\sigma})$ and $E_{\tilde{u}_{p}}(\pi^{\text{ref}},\tilde{\mu},\tilde{\sigma})$ lie within one standard deviation of the corresponding Monte Carlo error.
Comparing one test sample of these paths in Figure 6, we see that $\pi^{\text{NN}}_{\text{non rob}}$ follows a similar trajectory as $\pi^{\text{ref}}$, but it is shifted down by approximately $0.1$ in the market with $\tilde{\mu}=5.5\%,r=1.5\%$ (top, left sub-figure of Figure 6) and by approximately $0.05$ in the market with $\tilde{\mu}=10\%,r=3\%$ (bottom, left sub-figure of Figure 6). Moreover, sometimes when $\pi^{\text{ref}}$ hits the boundary of the no-trade region and subsequently stays there, $\pi^{\text{NN}}_{\text{non rob}}$ develops differently, as for instance for $t\approx(0.35,0.5)$ in the top, left sub-figure. Thus overall, even though we observed that the learned strategy $\pi^{\text{NN}}_{\text{non rob}}$ differs from $\pi^{\text{ref}}$, the resulting expected utilities in the reference market are nearly the same, i.e., $E_{\tilde{u}_{p}}(\pi^{\text{NN}}_{\text{non rob}},\tilde{\mu},\tilde{\sigma})\approx E_{\tilde{u}_{p}}(\pi^{\text{ref}},\tilde{\mu},\tilde{\sigma})$ in Table 3.
@@
However, when evaluated against the noisy market ($M_{\tilde{u}_{p}}$ entries of Table 3), $\pi^{\text{NN}}_{\text{non rob}}$ outperforms the reference strategy in the second digit. This seems reasonable, since we observed in Figure 6 how the optimized $\pi^{\text{NN}}_{\text{non rob}}$ tells to invest about 10% of wealth less into the risky asset than $\pi^{\text{ref}}$, and therefore acts similarly to a robustly trained strategy. Note that we observe such robustified behavior in $\pi^{\text{NN}}_{\text{non rob}}$ even though the training did not enforce it for $\pi^{\text{NN}}_{\text{non rob}}$. The robustly trained strategy $\pi^{\text{NN}}_{\text{fully rob}}$, however, outperforms both other strategies in the evaluation with noisy markets. Moreover, $\pi^{\text{NN}}_{\text{fully rob}}$ also beats the baseline $\pi^{\text{cash}}\equiv 0$ of investing everything in the risk-free asset, whose utility is given by $\tilde{u}_{p}(X_{0}\,e^{rT})$, in the noise-free and the noisy market evaluation. As we have seen before, this comes at the cost of smaller expected utility when evaluated in the reference market setting ($E_{\tilde{u}_{p}}$ entries of Table 3).
@@
This can also be seen in the histograms of the discounted terminal portfolio value $X_{T}$ in the noise-free reference market in Figure 7. Additionally, we see there that $\pi^{\text{NN}}_{\text{non rob}}$ has slightly smaller mean than $\pi^{\text{ref}}$, while also having smaller variance and being more right skewed. For the robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$ this is even more pronounced. In particular, $\pi^{\text{NN}}_{\text{non rob}}$ manages to achieve the same expected utility as the asymptotically optimal reference strategy $\pi^{\text{ref}}$, while substantially reducing the risk, measured for example by the discounted value at risk $\operatorname{VaR}_{\alpha}$ for $\alpha=5\%$ (Table 3). Indeed, with the initial wealth $X_{0}=1$, $\pi^{\text{NN}}_{\text{non rob}}$ reduces the value at risk compared to $\pi^{\text{ref}}$ by $6\%$ and $3\%$ in the small and large drift regimes, respectively. Moreover, the robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$ reduces the value at risk by $26\%$ and $21\%$, respectively.
@@
Test paths of $\pi^{\text{NN}}_{\text{fully rob}}$ are shown in the right sub-figures of Figure 6. Compared to their non-robust counterparts $\pi^{\text{NN}}_{\text{non rob}}$ in the left sub-figures, the proportional investment in the risky asset $\pi^{\text{NN}}_{\text{fully rob}}$ is held at much lower levels over the entire horizon.
@@
So far in this section, we found that across all considered market settings, NN-based non-robust strategies $\pi^{\text{NN}}_{\text{non rob}}$ that do not mimic asymptotically optimal investments $\pi^{\text{ref}}$ nonetheless perform on par with those asymptotically optimal reference strategies in terms of $E_{\tilde{u}_{p}}$. Moreover, we found that, interestingly, the training of NN-based strategies leads to reduced VaR of the corresponding terminal portfolio value distributions (cp. Table 3). For non-robust utility optimization in discrete time, real-world scenarios, this implies that there are suitable trading strategies beyond the classically known asymptotic strategy. Those alternative strategies potentially should be preferred if reducing VaR is of interest as well.

### v2

@@ 4.3.3 Evaluation Criterion: Early Stopping and General Assessment of Robustness
Since no explicit solution is known, we have to define new metrics to asses the robustness and quality of trained models, and to perform early stopping during training. We propose the following metrics, which are all based on a reference market model. In real world settings, this reference market model is obtained by selecting a period of historical test data on which the reference parameters $\tilde{\Sigma},\tilde{\mu}$ and their respective (co)variance are estimated. For synthetic experiments, we assume that reference parameters $\tilde{\Sigma},\tilde{\mu}$ and their (co)variance are given.
@@
    We remark that the cumulative noise type is the most realistic one for real world market behaviour, since parameter uncertainty increases with time. In the subsequent evaluations of Section˜4.3.4, we therefore focus on cumulative noise. In this setting, the market scenario $S^{j}$ is thus given by the SDE (12) with parameters $\Sigma^{\mathrm{noisy},j}_{t},\mu^{\mathrm{noisy},j}_{t}$ (where the Cholesky decomposition is used to compute a standard deviation matrix $\sigma$ from a variance matrix $\Sigma$).
@@ 4.3.4 Experimental Results
In this section, we report the results of a series of experiments on realistic penalized robust utility optimization as outlined earlier in Section˜4.3. In these experiments, we train and compare non-robust, robust volatility and fully robust models in a 2-dimensional market with setting (AS) for the reference volatility $\tilde{\Sigma}$, setting (AD) for the reference drift $\tilde{\mu}$, and all of interest rate, step size, and initial asset value as in Section˜4.1. We include proportional transaction costs $c_{\text{prop}}=1\%$, consider the three utility functions $u\in\{u_{1/2},u_{1}=\ln,u_{2}\}$ and path-dependent penalties as defined in Section˜4.3.2.
@@
For the model evaluation on the test set we use the standard deviations of $0.075$ for the volatility and $0.01$ for the drift. To counteract the too optimistic view of the validation metric (21), we double these standard deviations, i.e., $0.15$ for the volatility and $0.02$ for the drift, for the early stopping validation during training. Moreover, we use $n_{\text{pool}}=1000$, $B_{\text{test}}=B_{\text{val}}=40K$ test and validation samples and $160K$ training samples. As outlined earlier in Section˜4.3.3, we only report results with the cumulative noise type.
In Figure˜3 (left) we show the results of a grid search over the scaling factors $\lambda_{1},\lambda_{2}\in\{0.01,0.1,0.5,1,10,100\}$, which scale the penalties $R_{1}$ and $R_{2}$ for the quadratic covariation $l_{\text{QCV}}$ and for the average relative return $l_{\text{ARR}}$, respectively, for the 3 utility functions $u_{1/2},u_{1},u_{2}$.
@@
Moreover, in Table˜1 we see that for all 3 utility functions, the best combination of $\lambda_{1},\lambda_{2}$ leads to the pooled minimum expected utilities $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$, which are better (higher) than the baseline of investing everything in the risk-free bank account (given by $u(X_{0}\,e^{rT})$) and higher than the results of strategies trained in the non-robust setting, corresponding to $\lambda_{1}=\lambda_{2}=\infty$, as well as in the half-robust setting, where only the volatility is robustified, corresponding to $\lambda_{2}=\infty$ (here the maximum is always attained at $\lambda_{1}=0.01$).
@@
In particular, we see that the robust training leads to reasonable strategies, when using the appropriate penalty scaling parameters $\lambda_{1}$ and $\lambda_{2}$ for the given noisy evaluation market models and utility function. We also see that for all utility functions the optimal scaling parameters are close ($(\lambda_{1},\lambda_{2})=(0.5,10)$ for $u_{2}$, $(\lambda_{1},\lambda_{2})=(0.5,0.5)$ for $u_{1}$, $(\lambda_{1},\lambda_{2})=(1,1)$ for $u_{1/2}$). In real world settings this means that one has to first decide or understand against which type or magnitude of uncertainty one wants to robustify and then has to find appropriate penalty scaling factors for the respective setting, e.g., via grid search. Note that when comparing (minimal) expected utilities across different utility functions, it is important to keep in mind that larger utility values do not imply favorable distributions of the terminal wealth $X_{T}$, since different utility functions assign different values to the same terminal wealth (cf. Figure˜2).
In Figure˜4, we show test sample paths and mean $\pm$ standard deviations over all sample paths for the best strategy with utility $u_{1/2}$ evaluated against the trained market discriminator model and against one cumulative noisy market scenario. Especially in the average plots (the right panels of Figure˜4), we see that the trained market and the noisy market behave differently. Nevertheless, the trained strategy works well when evaluated against the noisy market model. In particular, the average behaviour of the strategy and the resulting average portfolio value evolution are very similar in both cases, even though the market behaves differently.
@@
Similarly to before in Sections˜4.1 and 4.2, FFNNs performed on par with RNNs (tested with the utility function $u_{1}$). This finding suggests that also in this more realistic and complex setup the optimal strategy is not path-dependent.
@@
In Figure˜3 (right) we show the results of a grid search over scaling factors $\lambda_{1},\lambda_{2}\in\{0.01,0.5,0.1,1.,10,100\}$ for the path-dependent penalties w.r.t. the quadratic covariation $l_{\text{QCV}}$ and the average relative return $l_{\text{ARR}}$, for the utility functions $u_{1/2},u_{1},$ and $u_{2}$. In Table˜2, we see that for all 3 utility functions, the best combination of $\lambda_{1},\lambda_{2}$ leads to the pooled-minimum expected utilities $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ that exceed the baseline of investing everything in the risk-free bank account (given by $u(X_{0}\,e^{rT})$). Moreover, $\max_{\lambda_{1},\lambda_{2}}M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ is also higher than the results of strategies trained in the non-robust setting, corresponding to $\lambda_{1}=\lambda_{2}=\infty$, as well as in the half-robust setting, where only the volatility is robustified, corresponding to $\lambda_{2}=\infty$ (which is always attained for $\lambda_{1}=0.01$).
In total, the evaluation against the noisy GARCH models shows that the robust model outperforms the others, when choosing the scaling factors correctly. Interestingly, the fully robust model for $u_{1}$ leads to a slightly higher value of $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ than for $u_{1/2}$. This is surprising, since $u_{1/2}$ dominates $u_{1}$ and one would expect this monotonicity to translate to the metric. [^4] In particular, using the strategy that results from the optimization w.r.t. $u_{1}$ in the evaluation with $u_{1/2}$ has to lead to the same or higher expected utility than the evaluation w.r.t. $u_{1}$. As we can see in Figure˜3, the best combination of scaling factors is always attained at a boundary point of the grid. Hence, it is expected that better scaling factors could be found, when further enlarging the grid. In line with this, we expect that then also the optimal combination for $u_{1/2}$ leads to a larger value of $M_{u}(\pi^{\text{NN}}_{\text{fully robust}})$ than for $u_{1}$.
@@
[…]
In Figure˜5 we compare the distribution of discounted terminal wealth $e^{-rT}\,X_{T}$ resulting from trading with the fully robust and non-robust strategies of Table˜1 $\pi^{\text{NN}}_{\text{fully rob}}$ and $\pi^{\text{NN}}_{\text{non rob}}$, respectively, obtained by (robust) portfolio optimization with the utility functions $u_{2}$ and $u_{1/2}$. We see that the trading with robust strategies leads to lower mean terminal portfolio value (PV), however, also manages to reduce the termianal PV’s standard deviation. Hence, even though there is some cost for the robustification, it also reduces the risk of the strategy. Additionally, the robust strategies have the advantage of being right-skewed (large profits are more likely to happen than large losses) compared to the non-robust ones which are not skewed (large profits and large losses are similarly likely to happen). Except for the skew, we see that changing from the non-robust (top-right) to the robust (top-left) strategy has a qualitatively similar effect as changing from a more risk-seeking (top-right) to a more risk-avers (bottom-right) utility function. Doing both (from top-right to bottom-left) increases the effect.
@@
In this section, we study in more detail the effect of small, proportional trading costs on optimal investments in both robust and non-robust utility optimization. In this paper, we so far tested how a GAN-based algorithm (cf. Section˜3) can be leveraged to recover known explicit solutions to robust utility optimization (Section˜4.1) and enable us to go beyond simple, friction-less settings to find robust strategies in presence of proportional trading costs and for general power utilities (Sections˜4.2 and 4.3). While we saw benefits of GAN-based trading strategies in expected utilities under noisy market evaluations in Sections˜4.2 and 4.3, in this Section˜4.4, we investigate potential benefits of the learned trading strategies compared to other well known, asymptotically optimal investments under vanishing trading costs.
In non-robust portfolio optimization, such an analysis of optimal trading strategies in presence of (vanishingly) small trading costs has been performed in Muhle-Karbe et al. (2017). For general power utilities of the form $\tilde{u}_{p}(x)=\tfrac{x^{1-p}}{1-p}$, for $p\in(0,1)$ (note that this definition slightly differs from our previous definition in Equation˜17), Muhle-Karbe et al. (2017) derive closed form approximations of asymptotically optimal trading strategies in presence of small trading costs for Black–Scholes and more general mean-reverting financial markets. In this section, we take up their discussion by comparing our GAN-based robust and non-robust strategies to the asymptotically optimal strategy derived in Muhle-Karbe et al. (2017) for Black–Scholes markets in Section˜4.4.1. Then, in Section˜4.4.2, we compare those strategies in one-dimensional markets consisting of a risky asset with heavy tailed log-returns.
@@ 4.4.1 Comparison in Black–Scholes Markets
In our experiments we use a similar setup as before with $p=0.5$, $c_{\text{prop}}=1\%$, $X_{0}=1$, $S_{0}=1$, $\tilde{\sigma}=35\%$, $\delta_{t}=1/65$ with $65$ steps, i.e., $T=1$, and two different regimes for the drift and risk-free rate $(\tilde{\mu}_{S},r)\in\{(4\%,1.5\%),(7\%,3\%)\}$, hence, $\tilde{\mu}\in\{5.5\%,10\%\}$. For the robust training we use the path-dependent penalties of Section˜4.3.2 and grid search over $\lambda_{1},\lambda_{2}\in\{0.01,0.5,0.1,1.,10,100\}$. For validation of robust models, we consider noisy market models around the reference market, with the same setup as in the corresponding paragraph in Section˜4.3.4. For the non-robust model, validation is done without noise, i.e., with the reference market with characteristics $\tilde{\mu}$ and $\tilde{\sigma}$ by approximating (13) with $B_{\text{val}}=40K$ independent validation sample paths. For each experiment, we run a range of models with various architectures and select the model that performed best in the validation for final evaluation on a separate test set. In the evaluation, we approximate the metric (20) with $n_{\text{pool}}=1000$ and $B_{\text{test}}=40K$, and the metric (13) with $B_{\text{test}}=100K$. Again, we did not see an advantage of using RNNs compared to standard FFNNs, therefore all results are reported for FFNNs.
@@ Results
In Table˜3, we show the comparison of the reference strategy $\pi^{\text{ref}}$ and the trained strategies $\pi^{\text{NN}}$ evaluated against the noisy market models and against the reference markets (with highest, i.e., best values per column in bold). For the fully robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$, we show the results of the best combination of scaling factors, given by $\lambda_{1}=10,\lambda_{2}=1$ (for $\tilde{\mu}=5.5\%$ and $r=1.5\%$), and by $\lambda_{1}=10,\lambda_{2}=10$ (for $\tilde{\mu}=10\%$ and $r=3\%$), respectively.
Table 3: Numerical results of reference strategy and trained GAN strategies evaluated against noisy market models with (20) and against the reference market with (13). Additionally, the discounted value at risk $\operatorname{VaR}_{\alpha}$ for $\alpha=5\%$ is computed in the noise-free reference markets, corresponding to the quantiles of the terminal portfolio value plotted in Figure˜7.
@@
[…]
@@
[…]
@@
However, when evaluated against the noisy market ($M_{\tilde{u}_{p}}$ entries of Table˜3), $\pi^{\text{NN}}_{\text{non rob}}$ outperforms the reference strategy in the second digit. This seems reasonable, since we observed in Figure˜6 how the optimized $\pi^{\text{NN}}_{\text{non rob}}$ tells to invest about 10% of wealth less into the risky asset than $\pi^{\text{ref}}$, and therefore acts similarly to a robustly trained strategy. Note that we observe such robustified behavior in $\pi^{\text{NN}}_{\text{non rob}}$ even though the training did not enforce it for $\pi^{\text{NN}}_{\text{non rob}}$. The robustly trained strategy $\pi^{\text{NN}}_{\text{fully rob}}$, however, outperforms both other strategies in the evaluation with noisy markets. Moreover, $\pi^{\text{NN}}_{\text{fully rob}}$ also beats the baseline $\pi^{\text{cash}}\equiv 0$ of investing everything in the risk-free asset, whose utility is given by $\tilde{u}_{p}(X_{0}\,e^{rT})$, in the noise-free and the noisy market evaluation. As we have seen before, this comes at the cost of smaller expected utility when evaluated in the reference market setting ($E_{\tilde{u}_{p}}$ entries of Table˜3).
@@
This can also be seen in the histograms of the discounted terminal portfolio value $X_{T}$ in the noise-free reference market in Figure˜7. Additionally, we see there that $\pi^{\text{NN}}_{\text{non rob}}$ has slightly smaller mean than $\pi^{\text{ref}}$, while also having smaller variance and being more right skewed. For the robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$ this is even more pronounced. In particular, $\pi^{\text{NN}}_{\text{non rob}}$ manages to achieve the same expected utility as the asymptotically optimal reference strategy $\pi^{\text{ref}}$, while substantially reducing the risk, measured for example by the discounted value at risk $\operatorname{VaR}_{\alpha}$ for $\alpha=5\%$ (Table˜3). Indeed, with the initial wealth $X_{0}=1$, $\pi^{\text{NN}}_{\text{non rob}}$ reduces the value at risk compared to $\pi^{\text{ref}}$ by $6\%$ and $3\%$ in the small and large drift regimes, respectively. Moreover, the robust strategy $\pi^{\text{NN}}_{\text{fully rob}}$ reduces the value at risk by $26\%$ and $21\%$, respectively.
@@
Test paths of $\pi^{\text{NN}}_{\text{fully rob}}$ are shown in the right sub-figures of Figure˜6. Compared to their non-robust counterparts $\pi^{\text{NN}}_{\text{non rob}}$ in the left sub-figures, the proportional investment in the risky asset $\pi^{\text{NN}}_{\text{fully rob}}$ is held at much lower levels over the entire horizon.
@@
So far in this section, we found that across all considered market settings, NN-based non-robust strategies $\pi^{\text{NN}}_{\text{non rob}}$ that do not mimic asymptotically optimal investments $\pi^{\text{ref}}$ nonetheless perform on par with those asymptotically optimal reference strategies in terms of $E_{\tilde{u}_{p}}$. Moreover, we found that, interestingly, the training of NN-based strategies leads to reduced VaR of the corresponding terminal portfolio value distributions (cp. Table˜3). For non-robust utility optimization in discrete time, real-world scenarios, this implies that there are suitable trading strategies beyond the classically known asymptotic strategy. Those alternative strategies should potentially be preferred if reducing VaR is of interest as well.